eviant:

Identify the vertex and the y-intercept of the graph of the function. y = 0.25(x + 5)2

eviant:

@Vocaloid

eviant:

@Tranquility

Vocaloid:

meant to get to this earlier but generally y + k = a(x-h)^2 gives a parabola with vertex (h,k) so compare this side by side to see what your h and k would be for the y-intercept just plug in x = 0 into y = 0.25(x + 5)2